OpenStudy (anonymous):

Sorry to bother but I need help thank you so very much. What is the first step in solving the problem, (3x-8)/(12)=9

OpenStudy (anonymous):

The first step is to turn \(\dfrac{3x-8}{12}\) into \(\dfrac{3x}{12}-\dfrac{8}{12}\).

OpenStudy (anonymous):

thank u very much

OpenStudy (precal):

or you can multiply both sides by 12

OpenStudy (anonymous):

Yeah, that's probably easier.

OpenStudy (precal):

 yes way easier